OpenStudy (anonymous):

what is the equation of the line of the line that has a slope of -1/3 and a y intercept of 5/2

OpenStudy (anonymous):

Slope of a line with Slope = m and y-intercept = c, is given by: \(y = mx + c\) Just plug in the values for m and c above, you will get your equation..

OpenStudy (anonymous):

-1/3x+5/2

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

You are missing the left hand side...

OpenStudy (anonymous):

\[y = -\frac{1}{3}x + \frac{5}{2}\]

OpenStudy (anonymous):

thankssss

OpenStudy (anonymous):

Sorry for coming late, I was busy at some other post...:)